A charge of 1.71 mC is placed at each corner of a square 0.170 m on a side.
Determine the magnitude of the force on each charge.
Determine the direction of the force on each charge.
Answer #1
direction of force along the diagonal line.

F = (9*10^9*1.71^2*10^-6/0.17^2)*(1.4142 + 0.5) = 1.746*10^6 N
Answer #2
\(2.273 \times 10^{6} N\)